# Augmenting Moment Retrieval: Zero-Dependency Two-Stage Learning

### v1

@@ 4.2 Implementation Details
Following previous methods [29, 59, 71, 69], we extract video features from the primary benchmark using CLIP [46] and SlowFast [11], while text features are extracted using the text branch of CLIP. For additional experiments on Charades-STA, we also employ VGG [57] or I3D [5] as video encoders and use GloVe [45] embeddings for text representation. Our training process consists of two stages, with 40 epochs in the first stage and 100 epochs in the second stage. In both stages, we use the AdamW optimizer with a learning rate of 1e-4 and a weight decay of 1e-4. The loss function weights are set as $\lambda_{\text{disc}}=0.5$ and $\lambda_{\text{dill}}=0.5$. All experiments are conducted on a single NVIDIA A40 GPU.
@@ 4.3 Comparison with State-of-the-arts
QVHighlights. As shown in Table 1, we compare our model with previous methods on both validation and test sets of QVHighlights. Compared to the previous state-of-the-art method, BAM-DETR [27], our model achieves significant improvements of +5.03% and +5.04% in R1@0.5 and R1@0.7 scores on the validation set, respectively, along with a +4.05% enhancement in average mAP. On the test set, our model maintains substantial advantages, notably surpassing BAM-DETR by +5.51% in R1@0.5.
@@
2D-TAN [79] VGG 40.94 22.85
UMT [34] VGG 48.31 29.25
MomentDiff [30] VGG 51.94 28.25
QD-DETR [41] VGG 52.77 31.13
TR-DETR [59] VGG 53.47 30.81
@@
VSLNet [76] I3D 47.31 30.19
QD-DETR [41] I3D 50.67 31.02
TR-DETR [59] I3D 55.51 33.66
@@
Charades-STA and TACoS. In Table 2, we further compare our model with previous methods on the Charades-STA and TACoS datasets. Additionally, Table 3 presents a comparison of different models on the Charades-STA dataset when replacing the backbone. Experimental results demonstrate that our model consistently achieves the best performance. Specifically, compared to the previous best-performing model UVCOM [69], our model achieves an average performance gain of +3.40% on Charades-STA and +3.64% on TACoS across different metrics. When replacing the backbone, our model also outperforms the previous state-of-the-art TR-DETR [59], with an average improvement of +4.51% for the VGG backbone and +3.27% for the I3D backbone across different metrics. These results highlight the strong robustness of our approach across various video types and feature extraction methods.
@@ 4.4 Ablation Study and Discussion
In this section, we conduct ablation studies on the validation set of QVHighlights to further analyze the effectiveness of our proposed components. Setting (a) represents our baseline model, which is the M-DETR [29] model combined with the local-global alignment module and visual feature refinement module proposed in [59]. In settings (b) and (c), we use data augmentation without the two-stage training strategy. Although the results of (c) are better than (b), both still underperform compared to the baseline model. This is due to the significant difference between the construction of augmented data and real data. If the model is directly trained on augmented data, it may overfit these synthetic samples and fail to adapt to the distribution of real data. Setting (e) shows another aspect; using only two-stage training without data augmentation (i.e., multi-round training on real data) does not improve model performance. However, when we combine data augmentation with two-stage training, model performance improves significantly, as shown in settings (e) and (f). This demonstrates that the two-stage training approach can leverage the advantages of augmented data and adapt to the distribution of real data. In settings (g) and (h), we further incorporate the dual-path distillation module into the two-stage training framework. This module can effectively distill knowledge from the first stage, further enhancing model performance. Settings (i) and (j) show the performance of adding the discriminative contrastive loss to the two-stage model. This loss improves the model’s semantic understanding ability by requiring it to distinguish ambiguous fragments. Finally, settings (k) and (l) use all modules of the two-stage training. Setting (l) also uses both data augmentation strategies simultaneously, achieving the best performance across all metrics.
@@ 4.5 Qualitative Results
As shown in Fig. 4, we conduct a qualitative comparison between our approach and recent open-source methods, UVCOM [69] and TaskWeave [71], on the validation split of QVHighlights. In the first case, the query text is “man in black shirt speaks in front of a tree at night.” In the video, immediately following the ground truth segment, another man wearing black clothing appears. However, he does not satisfy the condition “in front of a tree at night.” Both UVCOM and TaskWeave incorrectly localize this ambiguous segment, whereas our model successfully distinguishes it. In the second case, although all models identify the approximate location of the target segment, UVCOM and TaskWeave fail to precisely align its boundaries compared to our method. In the third case, while all models roughly identify the relevant segment, UVCOM and TaskWeave fail to accurately localize the boundaries between similar actions (e.g., throwing vs. rubbing), whereas our method achieves more precise segmentation. These results demonstrate the superiority of our model in semantic comprehension and precise localization.

### v2

@@ 4.2 Implementation Details
Following previous methods [lei2021qvhighlights, sun2024trdetr, yang2024taskweave, xiao2024uvcom], we extract video features from the primary benchmark using CLIP [radford2021clip] and SlowFast [feichtenhofer2019slowfast], while text features are extracted using the text branch of CLIP. For additional experiments on Charades-STA, we also employ VGG [simonyan2014vgg] or I3D [carreira2017i3d] as video encoders and use GloVe [pennington2014glove] embeddings for text representation. Our training process consists of two stages, with 40 epochs in the first stage and 100 epochs in the second stage. In both stages, we use the AdamW optimizer with a learning rate of 1e-4 and a weight decay of 1e-4. The loss function weights are set as $\lambda_{\text{disc}}=0.5$ and $\lambda_{\text{dill}}=0.5$. All experiments are conducted on a single NVIDIA A40 GPU.
@@ 4.3 Comparison with State-of-the-arts
QVHighlights. As shown in Table 1, we compare our model with previous methods on both validation and test sets of QVHighlights. Compared to the previous state-of-the-art method, BAM-DETR [lee2025bamdetr], our model achieves significant improvements of +5.03% and +5.04% in R1@0.5 and R1@0.7 scores on the validation set, respectively, along with a +4.05% enhancement in average mAP. On the test set, our model maintains substantial advantages, notably surpassing BAM-DETR by +5.51% in R1@0.5.
@@
2D-TAN [zhang2020tan] VGG 40.94 22.85
UMT [liu2022umt] VGG 48.31 29.25
MomentDiff [li2023momentdiff] VGG 51.94 28.25
QD-DETR [moon2023qd-detr] VGG 52.77 31.13
TR-DETR [sun2024trdetr] VGG 53.47 30.81
@@
VSLNet [zhang2020vslnet] I3D 47.31 30.19
QD-DETR [moon2023qd-detr] I3D 50.67 31.02
TR-DETR [sun2024trdetr] I3D 55.51 33.66
@@
Charades-STA and TACoS. In Table 2, we further compare our model with previous methods on the Charades-STA and TACoS datasets. Additionally, Table 3 presents a comparison of different models on the Charades-STA dataset when replacing the backbone. Experimental results demonstrate that our model consistently achieves the best performance. Specifically, compared to the previous best-performing model UVCOM [xiao2024uvcom], our model achieves an average performance gain of +3.40% on Charades-STA and +3.64% on TACoS across different metrics. When replacing the backbone, our model also outperforms the previous state-of-the-art TR-DETR [sun2024trdetr], with an average improvement of +4.51% for the VGG backbone and +3.27% for the I3D backbone across different metrics. These results highlight the strong robustness of our approach across various video types and feature extraction methods.
@@ 4.4 Ablation Study and Discussion
In this section, we conduct ablation studies on the validation set of QVHighlights to further analyze the effectiveness of our proposed components. Setting (a) represents our baseline model, which is the M-DETR [lei2021qvhighlights] model combined with the local-global alignment module and visual feature refinement module proposed in [sun2024trdetr]. In settings (b) and (c), we use data augmentation without the two-stage training strategy. Although the results of (c) are better than (b), both still underperform compared to the baseline model. This is due to the significant difference between the construction of augmented data and real data. If the model is directly trained on augmented data, it may overfit these synthetic samples and fail to adapt to the distribution of real data. Setting (e) shows another aspect; using only two-stage training without data augmentation (i.e., multi-round training on real data) does not improve model performance. However, when we combine data augmentation with two-stage training, model performance improves significantly, as shown in settings (e) and (f). This demonstrates that the two-stage training approach can leverage the advantages of augmented data and adapt to the distribution of real data. In settings (g) and (h), we further incorporate the dual-path distillation module into the two-stage training framework. This module can effectively distill knowledge from the first stage, further enhancing model performance. Settings (i) and (j) show the performance of adding the discriminative contrastive loss to the two-stage model. This loss improves the model’s semantic understanding ability by requiring it to distinguish ambiguous fragments. Finally, settings (k) and (l) use all modules of the two-stage training. Setting (l) also uses both data augmentation strategies simultaneously, achieving the best performance across all metrics.
@@ 4.5 Qualitative Results
As shown in Fig. 4, we conduct a qualitative comparison between our approach and recent open-source methods, UVCOM [xiao2024uvcom] and TaskWeave [yang2024taskweave], on the validation split of QVHighlights. In the first case, the query text is “man in black shirt speaks in front of a tree at night.” In the video, immediately following the ground truth segment, another man wearing black clothing appears. However, he does not satisfy the condition “in front of a tree at night.” Both UVCOM and TaskWeave incorrectly localize this ambiguous segment, whereas our model successfully distinguishes it. In the second case, although all models identify the approximate location of the target segment, UVCOM and TaskWeave fail to precisely align its boundaries compared to our method. In the third case, while all models roughly identify the relevant segment, UVCOM and TaskWeave fail to accurately localize the boundaries between similar actions (e.g., throwing vs. rubbing), whereas our method achieves more precise segmentation. These results demonstrate the superiority of our model in semantic comprehension and precise localization.